OpenStudy (anonymous):

-21(y-8)=y +(-67) Whats Y?

OpenStudy (anonymous):

Step 1 (distribute out all your parentheses): -21y + 168 = y - 67 Step 2 (put "like" terms on the same side): -21y - y = -67 -168 Step 3 (collect "like" terms): -22y = -235 Step 4 (get "y" by itself): y = (-235) / (-22) Therefore, y = 235/22 or in decimal form, y = 10.68